OpenStudy (anonymous):

Given the parent functions \[f(x)=3x+2\] and \[g(x)=5^x-10\] what is \[g(x)-f(x)\]

OpenStudy (campbell_st):

so you will have \[g(x) -f(x) = 5^x - 10 - (3x + 2)\] simplify and collect like terms

OpenStudy (anonymous):

Thank you so much. I kept coming up with "all real numbers" and that wasn't working for me...

OpenStudy (shamil98):

25 - 10 - 3x - 2 basically.

OpenStudy (anonymous):

I got \[-3x+5^x-12\] and that isn't one of my answer choices...

OpenStudy (campbell_st):

well i'd write it as \[g(x) - f(x) = 5^x - 3x - 12\]

OpenStudy (campbell_st):

but your answer is fine, you normally write with the highest power 1st.

OpenStudy (anonymous):

\[g(x)-f(x)=5^x-8-3x\] \[g(x)-f(x)=5^x-12-3x\] \[g(x)-f(x)=3x-8-5^x\] \[g(x)-f(x)=3x+12-5^x\]

OpenStudy (campbell_st):

ok... so you are looking for 5^x, -12 and -3x so which option...?

OpenStudy (anonymous):

The first one right?

OpenStudy (jdoe0001):

\(\bf g(x) - f(x) = 5^x - 12- 3x \implies g(x) - f(x) = 5^x - 3x - 12 \)

OpenStudy (jdoe0001):

the value and the signs are the same, the order doesn't quite matter, campbell_st just happen to have used the "standard form"

OpenStudy (anonymous):

So the second one?

OpenStudy (jdoe0001):

yes

OpenStudy (campbell_st):

thats correct. the 2nd one

OpenStudy (anonymous):

Thank you. Could you help me with another one like this?

OpenStudy (campbell_st):

if you post it in the normal manner there are lots of people who can help.